OpenStudy (anonymous):

Help needed in Related rates of change problem: Two quantities p and q are related by the equation (p-1)(q+2) = k where k is a constant. When p = 5 units, q is 7 and q is changing at the rate of 0.04 units per second. Find the rate at which point is changing. I've tried to work out dp/dt(what the question is asking) as (-36/25) *(q+2)^-2 but that is not what the exercise book is telling me. Can someone please clear this up?

OpenStudy (welshfella):

dp/dt = dq/dt * dp/dq = 0.04 * dp/dq

OpenStudy (welshfella):

Find the value of k:- (p-1)(q + 2) = k p = 5 when q = k so k = 4*9 = 36

OpenStudy (welshfella):

p-1)(q + 2) = 36 p- 1 = 36 / (q + 2) p = [36 / q + 2] + 1 Find dp/dq and substitute in the formula for dp/dt

OpenStudy (anonymous):

That's exactly what I did

OpenStudy (welshfella):

I don't understand how you got -36/25 Isnt it -36/81?

OpenStudy (welshfella):

(q + 2)^2 = (7 + 2)^2

OpenStudy (welshfella):

sorry but i have to go right now. You have the right method.

OpenStudy (anonymous):

dp/dt = dp/dq* dq/dt = -36(q+2)^-2 * 0.04 =(-36/25)*(q+2)^-2

OpenStudy (welshfella):

oh i see now all you need do is plug in q = 7 and you have your answer